OpenStudy (anonymous):

Solve the following system of equations. 2a – 3b + c = 10 2a – 2b – 2c = 2 a + 3b + 2c = –1

OpenStudy (ghazi):

use determinant to solve above equation

OpenStudy (anonymous):

idk wat that means

OpenStudy (amistre64):

there are different methods that all come to the same results; since you have not defined a method, the determinant is suitable

OpenStudy (anonymous):

the determinant is kidna useless you can tell if it's not factorable once you have the equation narrowed to 2 variables.

OpenStudy (anonymous):

just use substitution :D!

OpenStudy (amistre64):

ugh, substitution on these takes forever!!

OpenStudy (amistre64):

id eliminate :)

OpenStudy (anonymous):

how do u eliminatte

OpenStudy (amistre64):

you scale the eqauations so that you can combine them and eliminate one of more of the variables

OpenStudy (amistre64):

R1, 2a – 3b + c = 10 R2, 2a – 2b – 2c = 2 R3, a + 3b + 2c = –1 R3(-2), R2(-1) R1, 2a – 3b + c = 10 R2, -2a + 2b + 2c =-2 R3, -2a -6b -4c = 2 R1+R2, and R1+R3 R1, 2a – 3b + c = 10 R2, 0 - b + 3c = 8 R3, 0 - 9b - 3c = 12 R2(-9) R1, 2a – 3b + c = 10 R2, 0 +9b -27c = -72 R3, 0 - 9b - 3c = 12 R2 + R3 R1, 2a – 3b + c = 10 R2, 0 +9b -27c = -72 R3, 0 + 0 - 30c =-60

OpenStudy (amistre64):

-30c = -60 c = 2

OpenStudy (amistre64):

R1, 2a – 3b + (2) = 10 R2, 0 +9b -27(2) = -72 R1, 2a – 3b = 10-2 R2, 0 +9b = -72+54 b = -18/9 = -2 R1, 2a – 3(-2) = 8 R1, 2a + 6 = 8 a = (8-6)/2 = 1

OpenStudy (amistre64):

2a – 3b + c = 10 1 -2 2 ------------- 2 + 6 + 2 = 10 2a – 2b – 2c = 2 1 -2 2 -------------- 2 + 4 - 4 = 2 a + 3b + 2c = –1 1 -2 2 ------------------ 1 -6 + 4 = -1

OpenStudy (anonymous):

wow that is a lot of work i do not like these so much do you?

OpenStudy (amistre64):

i find them tedious

OpenStudy (amistre64):

it takes longer to show the work than it does to actually get the answer :)

OpenStudy (anonymous):

thankyou for ur help!

OpenStudy (amistre64):

i hope it was helpful :)